# Visual-Friendly Concept Protection via Selective Adversarial Perturbations

### v1

@@ 5. Experiments
5.1. Experimental Settings
Datasets. We used the CelebA-HQ (Karras et al., 2017) and VGGFace2 (Cao et al., 2018) datasets, following Anti-DreamBooth (Le et al., 2023) and MetaCloak (Liu et al., 2024), to validate the effectiveness of our method in privacy protection. CelebA-HQ is an enhanced version of the original CelebA dataset consisting of $30,000$ celebrity face images. VGGFace2 is a large-scale dataset with over $3.3$ million face images from $9,131$ unique identities. We randomly selected 50 identities from each dataset, ensuring that each had at least 15 images with a resolution exceeding 500 × 500 pixels. We also provide other non-face cases (artwork and landscape) in the paper.
Training details. The availability of Stable Diffusion pre-trained weights on Hugging Face  [^1] https://huggingface.co/ has significantly advanced research within the community. Here, our experiments primarily focus on the Stable Diffusion v1-4 version. For image pre-processing, we center crop and resize images to a resolution $512\times 512$. During protected image generation, we use $720$ iterations, step size $\alpha$ is $1/255$, and adversarial perturbation size $\epsilon$ is $8/255$. We test the protection performance on two types of SOTA personalization methods, i.e., Textual Inversion (TI) (Gal et al., 2022), and DreamBooth (DB) (Ruiz et al., 2023). During the training of the Textual Inversion, the constant learning rate is set to $5\times 10^{-4}$ with $3000$ optimization steps and batch size $1$. In DreamBooth, the constant learning rate is set to $5\times 10^{-7}$ with $1000$ optimization steps and batch size is $2$. We use a guidance scale of $7.5$ and $50$ denoising steps at the generation phase with a testing prompt. We set $c=0.1$ and $\alpha=0.5$ in practice. All experiments are complete on NVIDIA A100 GPU 40GB.
Baselines. We compare our method with five SOTA baselines, i.e., Mist (Liang et al., 2023), Anti-DreamBooth (referred to as Anti-DB) (Le et al., 2023), PhotoGuard (Salman et al., 2023), SDS(-) (Xue et al., 2024) and MetaCloak (Liu et al., 2024). In implementations, we set $\epsilon=8/255$ as the $\ell_{\infty}$ perturbation size, $\alpha=1/255$ as the step size for fair comparison.
  • Mist aims to maximize the training loss of Latent Diffusion Models (LDMs) while minimizing a textual loss between the protected image and a dummy image in the image encoder layer during protected image generation. In our implementation, following the original paper (Liang and Wu, 2023), we set the loss weight for the textual loss to $1\times 10^{-4}$ and run 720 iterations during protected image generation.
@@ 5.1. Experimental Settings
  • PhotoGuard targets diffusion-based inpainting/editing methods and offers two variants: one that attacks the VAE-encoder of LDMs, and another that attacks the final image outputs. We select the VAE-encoder variant due to its superior visual quality and set the training iterations to 200, as recommended in the original paper (Salman et al., 2023).
  • SDS(-) minimizes the SDS loss (Poole et al., 2022) to optimize adversarial data, achieving the best visual quality among all versions presented in its original paper (Xue et al., 2024). We set the training iterations to 100, following (Xue et al., 2024).
  • MetaCloak utilizes meta-learning to enhance the robustness of adversarial data against various transformations. We set the number of surrogate models to 5, the unrolling number to 1, the sample batch size to 1, and the training iterations to 4000, based on the recommendations in (Liu et al., 2024).
@@
Table 1. Automatic and human evaluations across different methods, including their respective mean and standard deviation ($\pm$) on the VGGFace2 and CelebA-HQ datasets. All methods employed a perturbation size of $\epsilon$ = 8/255. In the HumanEval results, “Loss/Tie/Win” in “Visibility” indicates the voting rates for VCPro losses, ties, and wins in comparison to other methods regarding the stealthiness of their adversarial perturbations. “TI/DB” represents the percentage of images generated by Textual Inversion/DreamBooth identified as synthetic. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
VGGFace2 No Defense - - - - $3.61_{\pm 0.98}$ $0.24_{\pm 0.34}$ $3.95_{\pm 1.11}$ $0.38_{\pm 0.25}$ 64 : 26 : 10 - -
Mist arXiv23 $105.93_{\pm 9.76\hphantom{0}}$ $0.83_{\pm 0.03}$ $32.48_{\pm 1.12}$ $1.24_{\pm 0.41}$ $0.11_{\pm 0.29}$ $1.02_{\pm 0.11}$ $0.25_{\pm 0.28}$ 04 : 08 : 88 098.65 097.26
Anti-DreamBooth ICCV23 $\hphantom{0}96.24_{\pm 10.44}$ $0.82_{\pm 0.03}$ $32.43_{\pm 1.10}$ $1.18_{\pm 0.36}$ $0.05_{\pm 0.29}$ $1.03_{\pm 0.14}$ $0.26_{\pm 0.28}$ 03 : 09 : 88 097.33 100.00
PhotoGuard ICML23 $\hphantom{0}62.07_{\pm 8.15\hphantom{0}}$ $0.84_{\pm 0.03}$ $33.17_{\pm 1.21}$ $1.50_{\pm 0.54}$ $0.13_{\pm 0.34}$ $1.26_{\pm 0.30}$ $0.27_{\pm 0.31}$ 16 : 21 : 63 098.72 083.95
SDS(-) ICLR24 $\hphantom{0}47.48_{\pm 6.81\hphantom{0}}$ $0.81_{\pm 0.03}$ $32.91_{\pm 0.43}$ $2.09_{\pm 0.82}$ $0.07_{\pm 0.37}$ $2.74_{\pm 0.97}$ $0.33_{\pm 0.26}$ 14 : 17 : 69 094.44 077.03
MetaCloak CVPR24 $204.25_{\pm 54.52}$ $0.82_{\pm 0.02}$ $32.09_{\pm 0.63}$ $1.86_{\pm 0.62}$ $0.19_{\pm 0.33}$ $1.33_{\pm 0.41}$ $0.25_{\pm 0.32}$ 02 : 13 : 85 091.18 090.41
VCPro $\hphantom{0}27.04_{\pm 4.38\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.33_{\pm 1.79}$ $2.31_{\pm 0.78}$ $0.03_{\pm 0.32}$ $2.06_{\pm 0.70}$ $0.21_{\pm 0.34}$ 098.57 097.37
CelebA-HQ No Defense - - - $4.40_{\pm 0.86}$ $0.41_{\pm 0.22}$ $4.84_{\pm 0.40}$ $0.63_{\pm 0.13}$ 79 : 18 : 03 - -
Mist arXiv23 $\hphantom{0}78.32_{\pm 11.83}$ $0.86_{\pm 0.03}$ $33.75_{\pm 0.21}$ $1.45_{\pm 0.58}$ $0.26_{\pm 0.26}$ $1.04_{\pm 0.11}$ $0.44_{\pm 0.21}$ 06 : 03 : 91 096.20 100.00
Anti-DreamBooth ICCV23 $\hphantom{0}78.45_{\pm 11.66}$ $0.86_{\pm 0.03}$ $33.74_{\pm 0.21}$ $1.81_{\pm 1.20}$ $0.28_{\pm 0.25}$ $1.04_{\pm 0.15}$ $0.44_{\pm 0.21}$ 00 : 05 : 95 096.92 096.51
PhotoGuard ICML23 $\hphantom{0}45.18_{\pm 7.21\hphantom{0}}$ $0.88_{\pm 0.02}$ $35.08_{\pm 0.25}$ $1.81_{\pm 0.59}$ $0.30_{\pm 0.24}$ $1.23_{\pm 0.35}$ $0.48_{\pm 0.17}$ 04 : 21 : 75 094.12 094.37
SDS(-) ICLR24 $\hphantom{0}34.53_{\pm 6.16\hphantom{0}}$ $0.82_{\pm 0.03}$ $33.01_{\pm 0.28}$ $2.29_{\pm 0.70}$ $0.09_{\pm 0.28}$ $2.92_{\pm 0.74}$ $0.55_{\pm 0.15}$ 09 : 10 : 81 098.70 084.51
MetaCloak CVPR24 $161.93_{\pm 26.93}$ $0.86_{\pm 0.03}$ $33.12_{\pm 0.20}$ $1.47_{\pm 0.56}$ $0.25_{\pm 0.26}$ $1.59_{\pm 0.53}$ $0.44_{\pm 0.16}$ 00 : 05 : 95 096.00 095.52
VCPro $\hphantom{0}16.21_{\pm 2.89\hphantom{0}}$ $0.95_{\pm 0.01}$ $39.32_{\pm 0.40}$ $2.61_{\pm 1.00}$ $0.09_{\pm 0.28}$ $2.62_{\pm 0.95}$ $0.44_{\pm 0.19}$ 100.00 095.64
@@
Figure 3. Qualitative defense results for both the human face and non-face cases. For generating protected images, we use an $\epsilon$ value of 8/255 with Stable Diffusion v1-4. Each row displays the original images (No Defense), the protected images generated by PhotoGuard, Anti-DreamBooth, Mist, SDS(-), MetaCloak, and our method, along with their respective outcomes in Textual Inversion (TI) and DreamBooth (DB). Additionally, we include visualizations of adversarial perturbations represented as black-yellow images. For better detail, please zoom in.
5.2. Main Results
Quantitative Results.
Table 1 shows the automatic and human evaluation results for the visibility of protected images and the final protection effectiveness of different methods.
@@ 5.2. Main Results
Qualititative Results. Fig. 3 illustrates protected images and protective effects on individual identity information and landscapes of different methods under Textual Inversion and DreamBooth. We also visualize adversarial perturbations by regularizing all perturbations to the range [0,1] and then visualizing using colormap.
@@
Under the protection of our method, the personalization methods fail to generate visual-fidelity results. As shown in the case of important information such as faces, landmark landscapes, or buildings, our adversarial perturbation is less obvious (especially in the background part), and effectively distorts textures and features in important regions, thereby preventing the malicious personalization of individual identification information or copyright, so that the adversarial perturbation is more subtle while achieving the protection purpose, and find a better balance between the protection effect and the perturbation intensity, which reduces user concerns about image quality degradation after protection. PhotoGuard, Mist, Anti-DreamBooth, and MetaCloak tend to add strange textures to the whole image and show obvious adversarial perturbations (the second, third, fourth, and sixth columns). SDS(-) can cause generated images to be too blurry and has limited protection for the DreamBooth. Its protected image is like adding strange circular blobs to the original image (the fifth column). Visualization of its perturbations also supports the idea that excessive brightness leads to lower SSIM values. These approaches serve as a protective measure, ensuring that certain image regions exhibit unique textures. The presence of such textures can strongly suggest infringement. We also find that the protection difficulty of Textual Inversion is far less than that of DreamBooth because DreamBooth usually fine-tunes most parameters in Stable Diffusion. At the same time, Textual Inversion only adds a few new word embeddings. It is also supported in Table 1.
@@
User Study. We evaluate the performance of VCPro in the concept protection task with five selected methods, i.e., PhotoGuard (Salman et al., 2023), Mist (Liang and Wu, 2023), Anti-DreamBooth (Le et al., 2023), SDS(-)(Xue et al., 2024), and MetaCloak (Liu et al., 2024). A total of 27 participants took part in the survey, and they all use social media proficiently. We randomly sampled protected images and generation images from all 100 identities in VGGFace2 and CelebA-HQ and six groups of non-face data. The study consists of three parts, including visibility of adversarial perturbations (1,116 valid votes), protection effects under Textual Inversion (1,078 valid votes), and protection effects under DreamBooth (1,067 valid votes).
User Study I
@@
In the survey on the visibility of adversarial perturbations, participants were shown a VCPro-protected image alongside a protected image generated by another method based on the same original image. They were then asked to vote on the visual quality, choosing between ”A wins,” ”tie,” or ”B wins.” The visual quality of VCPro was significantly better than that of other methods. Specifically, when compared to Mist, VCPro received 88.27% of the votes (6.79% ties and 4.94% for Mist); against Anti-DreamBooth, VCPro received 91.72% (7.01% ties and 1.27% for Anti-DreamBooth); against MetaCloak, VCPro received 90.68% (8.70% ties and 0.62% for MetaCloak); against PhotoGuard, VCPro received 70.25% (20.89% ties and 8.86% for PhotoGuard); and against SDS(-), VCPro received 76.97% (12.73% ties and 10.30% for SDS(-)).
User Study II
@@
In the survey on protection efficacy under Textual Inversion, participants were given an original image $x$ and asked to determine whether the image generated through Textual Inversion, trained on protected image $x^{\prime}$ from different methods, was synthetic. The results were as follows: for Mist, 98.58% voted synthetic; for Anti-DreamBooth, 96.08%; for MetaCloak, 94.19%; for PhotoGuard, 96.79%; for SDS(-), 96.08%; and for VCPro, 98.66%.
User Study III
In the survey on protection efficacy under DreamBooth, participants were shown an original image and a DreamBooth-generated image trained on its protected images from different methods. They were then asked to determine whether the latter was synthetic. The results were as follows: for Mist, 98.50% voted synthetic; for Anti-DreamBooth, 98.00%; for MetaCloak, 92.86%; for PhotoGuard, 86.93%; for SDS(-), 83.55%; and for VCPro, 95.21%. Most participants also acknowledged that VCPro had a protective effect.
[…]
5.3. Perturbation Size Influence
Perturbation size $\epsilon$ controls the maximum allowable change in pixel values of adversarial perturbation. As shown in Fig. 4 and Table 2, different levels of adversarial perturbation size noticeably influence the protection outcomes: Compared to low $\epsilon$, large $\epsilon$ presents worse invisibility of adversarial perturbations while the more obvious protection effects.
Table 2. The influence of different perturbation size $\epsilon$ on VGGFace2 datasets. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Budget Visibility of Perturbations Textual Inversion DreamBooth
0 - - - 3.61±0.98 00.24±0.34 3.95±1.11 0.38±0.25
@@ 5.3. Perturbation Size Influence
Figure 5. The ablation experiments of different mask sizes and modules. Anti-DreamBooth (Anti-DB) is our baseline. The red region indicates the important region. Here Input-mask means optimizing adversarial perturbations by directly restricting the mask range. Zoom in for a better view.
5.4. Ablation Study
As shown in Fig. 5 and Table 3, we conducted ablation experiments of two modules in VCPro: Regional Adversarial Loss (RAL) in Sec. 4.2 and Lagrangian Multiplier-based Loose Solution (LMLS) in Sec. 4.3.
Compared with our baseline Anti-DreamBooth, we find that Anti-DreamBooth+RAL reduces the visibility of the perturbations, while the protection effect is concentrated in the mask area, resulting in a slight decrease in the global image quality assessment LIQE but stable in CLIP-FACE. At the same time, we also consider the effect of different mask ranges, smaller masks will further reduce the visibility of the perturbations, and the protection range is further reduced.
LMLS shows a similar phenomenon, where Anti-DreamBooth+LMLS reduces the visibility of the noise and tends to leave the high-frequency part, while the final protection effect also leaves a streaky texture across the whole image. That is to say, the adversarial perturbation added to a specific space and frequency domain will directly affect the protection effect in the corresponding space and frequency domains.
Moreover, only push-loss without pull-loss in RAL (VCPro (w/o Pull) in Fig. 5 and Table 3) will reduce the training loss gradient feedback about the protection effect, while the loss that constrains the visibility of the perturbations is not affected resulting in decreased visibility of the perturbations and the protection effect.
@@ 5.4. Ablation Study
Table 3. Automatic evaluation results on ablation experiments with the corresponding mean and standard deviation ($\pm$). All methods use $\epsilon$ = 8/255 on VGGFace2 dataset. ↑ means the higher the better, and vice versa. The best result in each column is in bold. Here Input-Mask is to constrain the adversarial perturbations directly on the input image to be only within the mask, we are constraining on the optimization objective, and the adversarial perturbations will be present in the whole image.
@@
Sensitivity Analysis of Training Interation

### v2

@@ V Experiments
V-A Experimental Settings
Datasets. We used the CelebA-HQ [34] and VGGFace2 [35] datasets, following Anti-DreamBooth [8] and MetaCloak [10], to validate the effectiveness of our method in privacy protection. CelebA-HQ is an enhanced version of the original CelebA dataset consisting of $30,000$ celebrity face images. VGGFace2 is a large-scale dataset with over $3.3$ million face images from $9,131$ unique identities. We randomly selected 50 identities from each dataset, ensuring that each had at least 15 images with a resolution exceeding 500 × 500 pixels. We also provide other non-face cases (artwork and landscape) in the paper.
Training details. The availability of Stable Diffusion pre-trained weights on Hugging Face  [^1] https://huggingface.co/ has significantly advanced research within the community. Here, our experiments primarily focus on the Stable Diffusion v1-4 version. For image pre-processing, we center crop and resize images to a resolution $512\times 512$. During protected image generation, we use $720$ iterations, step size $\alpha$ is $1/255$, and adversarial perturbation size $\epsilon$ is $8/255$. We test the protection performance on two types of SOTA personalization methods, i.e., Textual Inversion (TI) [6], and DreamBooth (DB) [7]. During the training of the Textual Inversion, the constant learning rate is set to $5\times 10^{-4}$ with $3000$ optimization steps and batch size $1$. In DreamBooth, the constant learning rate is set to $5\times 10^{-7}$ with $1000$ optimization steps and batch size is $2$. We use a guidance scale of $7.5$ and $50$ denoising steps at the generation phase with a testing prompt. We set $c=0.1$ and $\alpha=0.5$ in practice. All experiments are complete on NVIDIA A100 GPU 40GB.
Baselines. We compare our method with five SOTA baselines, i.e., Mist [13], Anti-DreamBooth (referred to as Anti-DB) [8], PhotoGuard [12], SDS(-) [11] and MetaCloak [10]. In implementations, we set $\epsilon=8/255$ as the $\ell_{\infty}$ perturbation size, $\alpha=1/255$ as the step size for fair comparison.
  • Mist aims to maximize the training loss of Latent Diffusion Models (LDMs) while minimizing a textual loss between the protected image and a dummy image in the image encoder layer during protected image generation. In our implementation, following the original paper [14], we set the loss weight for the textual loss to $1\times 10^{-4}$ and run 720 iterations during protected image generation.
@@ V-A Experimental Settings
  • PhotoGuard targets diffusion-based inpainting/editing methods and offers two variants: one that attacks the VAE-encoder of LDMs, and another that attacks the final image outputs. We select the VAE-encoder variant due to its superior visual quality and set the training iterations to 200, as recommended in the original paper [12].
  • SDS(-) minimizes the SDS loss [36] to optimize adversarial data, achieving the best visual quality among all versions presented in its original paper [11]. We set the training iterations to 100, following [11].
  • MetaCloak utilizes meta-learning to enhance the robustness of adversarial data against various transformations. We set the number of surrogate models to 5, the unrolling number to 1, the sample batch size to 1, and the training iterations to 4000, based on the recommendations in [10].
@@
TABLE I: Automatic and human evaluations across different methods, including their respective mean and standard deviation ($\pm$) on the VGGFace2 and CelebA-HQ datasets. All methods employed a perturbation size of $\epsilon$ = 8/255. In the HumanEval results, “Loss/Tie/Win” in “Visibility” indicates the voting rates for VCPro losses, ties, and wins in comparison to other methods regarding the stealthiness of their adversarial perturbations. “TI/DB” represents the percentage of images generated by Textual Inversion/DreamBooth identified as synthetic. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
VGGFace2 No Defense - - - - $3.61_{\pm 0.98}$ $0.24_{\pm 0.34}$ $3.95_{\pm 1.11}$ $0.38_{\pm 0.25}$ 64 : 25 : 11 - -
Mist arXiv23 $105.93_{\pm 9.76\hphantom{0}}$ $0.83_{\pm 0.03}$ $32.48_{\pm 1.12}$ $1.24_{\pm 0.41}$ $0.11_{\pm 0.29}$ $1.02_{\pm 0.11}$ $0.25_{\pm 0.28}$ 04 : 06 : 90 098.34 097.23
Anti-DreamBooth ICCV23 $\hphantom{0}96.24_{\pm 10.44}$ $0.82_{\pm 0.03}$ $32.43_{\pm 1.10}$ $1.18_{\pm 0.36}$ $0.05_{\pm 0.29}$ $1.03_{\pm 0.14}$ $0.26_{\pm 0.28}$ 03 : 09 : 88 097.38 100.00
PhotoGuard ICML23 $\hphantom{0}62.07_{\pm 8.15\hphantom{0}}$ $0.84_{\pm 0.03}$ $33.17_{\pm 1.21}$ $1.50_{\pm 0.54}$ $0.13_{\pm 0.34}$ $1.26_{\pm 0.30}$ $0.27_{\pm 0.31}$ 13 : 19 : 68 098.70 084.02
SDS(-) ICLR24 $\hphantom{0}47.48_{\pm 6.81\hphantom{0}}$ $0.81_{\pm 0.03}$ $32.91_{\pm 0.43}$ $2.09_{\pm 0.82}$ $0.07_{\pm 0.37}$ $2.74_{\pm 0.97}$ $0.33_{\pm 0.26}$ 13 : 16 : 71 094.56 77.13
MetaCloak CVPR24 $204.25_{\pm 54.52}$ $0.82_{\pm 0.02}$ $32.09_{\pm 0.63}$ $1.86_{\pm 0.62}$ $0.19_{\pm 0.33}$ $1.33_{\pm 0.41}$ $0.25_{\pm 0.32}$ 02 : 10 : 88 091.03 090.44
VCPro $\hphantom{0}27.04_{\pm 4.38\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.33_{\pm 1.79}$ $2.31_{\pm 0.78}$ $0.03_{\pm 0.32}$ $2.06_{\pm 0.70}$ $0.21_{\pm 0.34}$ 098.61 097.35
CelebA-HQ No Defense - - - $4.40_{\pm 0.86}$ $0.41_{\pm 0.22}$ $4.84_{\pm 0.40}$ $0.63_{\pm 0.13}$ 77 : 19 : 04 - -
Mist arXiv23 $\hphantom{0}78.32_{\pm 11.83}$ $0.86_{\pm 0.03}$ $33.75_{\pm 0.21}$ $1.45_{\pm 0.58}$ $0.26_{\pm 0.26}$ $1.04_{\pm 0.11}$ $0.44_{\pm 0.21}$ 05 : 04 : 91 096.29 99.99
Anti-DreamBooth ICCV23 $\hphantom{0}78.45_{\pm 11.66}$ $0.86_{\pm 0.03}$ $33.74_{\pm 0.21}$ $1.81_{\pm 1.20}$ $0.28_{\pm 0.25}$ $1.04_{\pm 0.15}$ $0.44_{\pm 0.21}$ 00 : 04 : 96 097.01 096.53
PhotoGuard ICML23 $\hphantom{0}45.18_{\pm 7.21\hphantom{0}}$ $0.88_{\pm 0.02}$ $35.08_{\pm 0.25}$ $1.81_{\pm 0.59}$ $0.30_{\pm 0.24}$ $1.23_{\pm 0.35}$ $0.48_{\pm 0.17}$ 02 : 22 : 76 094.25 094.31
SDS(-) ICLR24 $\hphantom{0}34.53_{\pm 6.16\hphantom{0}}$ $0.82_{\pm 0.03}$ $33.01_{\pm 0.28}$ $2.29_{\pm 0.70}$ $0.09_{\pm 0.28}$ $2.92_{\pm 0.74}$ $0.55_{\pm 0.15}$ 09 : 11 : 80 098.63 084.32
MetaCloak CVPR24 $161.93_{\pm 26.93}$ $0.86_{\pm 0.03}$ $33.12_{\pm 0.20}$ $1.47_{\pm 0.56}$ $0.25_{\pm 0.26}$ $1.59_{\pm 0.53}$ $0.44_{\pm 0.16}$ 00 : 04 : 96 096.02 095.48
VCPro $\hphantom{0}16.21_{\pm 2.89\hphantom{0}}$ $0.95_{\pm 0.01}$ $39.32_{\pm 0.40}$ $2.61_{\pm 1.00}$ $0.09_{\pm 0.28}$ $2.62_{\pm 0.95}$ $0.44_{\pm 0.19}$ 100.00 095.67
@@
Figure 3: Qualitative defense results for both the human face and non-face cases. For generating protected images, we use an $\epsilon$ value of 8/255 with Stable Diffusion v1-4. Each row displays the original images (No Defense), the protected images generated by PhotoGuard, Anti-DreamBooth, Mist, SDS(-), MetaCloak, and our method, along with their respective outcomes in Textual Inversion (TI) and DreamBooth (DB). Additionally, we include visualizations of adversarial perturbations represented as black-yellow images. For better detail, please zoom in.
V-B Main Results
Quantitative Results. Table I shows the automatic and human evaluation results for the visibility of protected images and the final protection effectiveness of different methods.
@@ V-B Main Results
Qualititative Results. Fig. 3 illustrates protected images and protective effects on individual identity information and landscapes of different methods under Textual Inversion and DreamBooth. We also visualize adversarial perturbations by regularizing all perturbations to the range [0,1] and then visualizing using a specific colormap.
@@
Under the protection of our method, the personalization methods fail to generate visual-fidelity results. As shown in the case of important information such as faces, landmark landscapes, or buildings, our adversarial perturbation is less noticeable, especially in the background, and effectively distorts textures and features in crucial areas. This prevents the unauthorized use of personal identification information or copyrighted material, making the perturbation more subtle while still serving its protective purpose. This approach strikes a better balance between safeguarding the content and maintaining its quality, thereby reducing user concerns about image degradation after protection. PhotoGuard, Mist, Anti-DreamBooth, and MetaCloak tend to add strange textures to the whole image and show obvious adversarial perturbations (the second, third, fourth, and sixth columns). SDS(-) can cause generated images to be too blurry and has limited protection for the DreamBooth. Its protected image is like adding strange circular blobs to the original image (the fifth column). Visualization of its perturbations also supports the idea that excessive brightness leads to lower SSIM values. These approaches serve as a protective measure, ensuring that certain image regions exhibit unique textures. The presence of such textures can strongly suggest infringement. We also find that the protection difficulty of Textual Inversion is far less than that of DreamBooth because DreamBooth usually fine-tunes most parameters in Stable Diffusion. At the same time, Textual Inversion only adds a few new word embeddings. It is also supported in Table I.
User Study. We evaluate the performance of VCPro in the concept protection task with five selected methods, i.e., PhotoGuard [12], Mist [14], Anti-DreamBooth [8], SDS(-)[11], and MetaCloak [10]. A total of 50 participants took part in the survey, and they all use social media proficiently. Participants comprise 58% male and 42% female, with ages ranging from 18 to 55 years (mean: 24.5). All participants have normal color vision and social media proficiency relevant to our image protection context. Educational backgrounds span from undergraduate to doctoral level (undergraduate: 54% graduate: 26%, doctoral: 4%). Professional expertise was distributed across computer science (50%), electronic information technology (36%), and visual arts (14%). For quality control, all participants completed preliminary training and visual acuity testing. We implemented consistency checks through randomized question repetition with shuffled options, requiring participants to maintain at least 95% response consistency for inclusion in the final analysis.
@@
We randomly sampled protected images and generated images from all 100 identities in VGGFace2 and CelebA-HQ and six groups of non-face data. The study consists of three parts, including visibility of adversarial perturbations (2064 valid votes), protection effects under Textual Inversion (1996 valid votes), and protection effects under DreamBooth (1990 valid votes).
@@
User Study I. In the survey on the visibility of adversarial perturbations, participants were shown a VCPro-protected image alongside a protected image generated by another method based on the same original image. They were then asked to vote on the visual quality, choosing between “A wins”, “tie”, or “B wins”. The visual quality of VCPro was significantly better than that of other methods. As shown in Table I, our VCPro significantly outperformed all baseline methods in direct comparisons. VCPro received between 68% and 96% of participant votes when compared against competing methods. To assess the reliability of our user study results, we calculated the Kendall coefficient [41] as 0.71 ($p<0.05$), indicating substantial inter-rater agreement among participants.
@@
User Study II. In the survey on protection efficacy under Textual Inversion, participants were given an original image $x$ and asked to determine whether the image generated through Textual Inversion, trained on protected image $x^{\prime}$ from different methods, was synthetic. As shown in Table I, all protection methods received at least 91% of synthetic votes, with VCPro demonstrating the most effective protection as acknowledged by the majority of participants. To evaluate the consistency of the user study, we calculated Cohen’s Kappa [42], which yielded a value of 0.88 ($p<0.05$), indicating a strong agreement among participants.
User Study III. In the survey on protection efficacy under DreamBooth, as shown in Table 1, participants recognized an average 97.01% protective success rate of VCPro. To measure the consistency of the study, we calculated Cohen’s Kappa, which resulted in a value of 0.88 ($p<0.05$).
@@
Figure 4: The ablation experiments of different mask sizes and modules. Anti-DreamBooth (Anti-DB) is our baseline. The red region indicates the important region. Here Input-mask means optimizing adversarial perturbations by directly restricting the mask range. Zoom in for a better view.
@@
TABLE II: Automatic evaluation results on ablation experiments with the corresponding mean and standard deviation ($\pm$). All methods use $\epsilon$ = 8/255 on VGGFace2 dataset. ↑ means the higher the better, and vice versa. The best result in each column is in bold. Here Input-Mask is to constrain the adversarial perturbations directly on the input image to be only within the mask, we are constraining on the optimization objective, and the adversarial perturbations will be present in the whole image.
@@
[…]
V-C Ablation Study
As shown in Fig. 4 and Table II, we conducted ablation experiments of two modules in VCPro: Regional Adversarial Loss (RAL) in Sec. IV-B and Lagrangian Multiplier-based Loose Solution (LMLS) in Sec. IV-C. Anti-DreamBooth is our baseline.
Compared with Anti-DreamBooth, we find that Anti-DreamBooth+RAL reduces the visibility of the perturbations, while the protection effect is concentrated in the mask area, resulting in a slight decrease in the global image quality assessment LIQE but stable in CLIP-FACE. At the same time, we also consider the effect of different mask ranges, smaller masks will further reduce the visibility of the perturbations, and the protection range is further reduced.
LMLS shows a similar phenomenon where Anti-DreamBooth+LMLS reduces the visibility of the noise and tends to leave the high-frequency part, while the final protection effect also leaves a streaky texture across the whole image. Consequently, the adversarial perturbation added to a specific space and frequency domain will directly affect the protection effect in the corresponding space and frequency domains.
Moreover, only push-loss without pull-loss in RAL (VCPro (w/o Pull) in Fig. 4 and Table II) will reduce the training loss gradient feedback about the protection effect, while the loss that constrains the visibility of the perturbations is not affected resulting in decreased visibility of the perturbations and the protection effect.
V-D Perturbation Size Influence
Perturbation size $\epsilon$ controls the maximum allowable change in pixel values of adversarial perturbation. As shown in Fig. 5 and Table III, different levels of adversarial perturbation size noticeably influence the protection outcomes: Compared to low $\epsilon$, large $\epsilon$ presents worse invisibility of adversarial perturbations while the more obvious protection effects.
TABLE III: The influence of different perturbation size $\epsilon$ on VGGFace2 datasets. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Budget Visibility of Perturbations Textual Inversion DreamBooth
0 - - - 3.61±0.98 00.24±0.34 3.95±1.11 0.38±0.25
V-E Hyper-parameter Analysis
Here we provide quantitative results to study the influence of hyper-parameters: training iteration, $c$, and $\alpha$. We also provide qualitative results in supplementary materials.